OpenStudy (anonymous):

what is the sum of the arithmetic sequence 135,123,111... if there are 34 terms?

OpenStudy (amistre64):

what is the difference between the 2nd term and the 1st term?

OpenStudy (amistre64):

123 - 135 = ?

OpenStudy (anonymous):

12.

OpenStudy (anonymous):

Dude ,Try to solve your own exam by your own ,believe me.....u gonna solve it faster.

OpenStudy (amistre64):

lets go with -12 :) this is our common difference (d).

OpenStudy (anonymous):

ok.

OpenStudy (amistre64):

we can use the formula: an = a1 + d(n-1) to find the last term that we will need

OpenStudy (amistre64):

when n=34 we get: a34 = 135 -12(33) whats tat get us?

OpenStudy (anonymous):

a34= 135- 396?

OpenStudy (amistre64):

keep going, so far so good

OpenStudy (anonymous):

i dont kow how to keep going.

OpenStudy (amistre64):

135 - 396 = ??

OpenStudy (anonymous):

-261

OpenStudy (amistre64):

good; this is our last term in the sequece that we need

OpenStudy (amistre64):

now we need to use another nifty little formula: \[Sn=\frac{n}{2}(a1+an)\]

OpenStudy (amistre64):

\[S_{34}=\frac{34}{2}(135-261)\]

OpenStudy (anonymous):

@amistre64 where can I find proof for that formula?

OpenStudy (anonymous):

okay... now?

OpenStudy (amistre64):

now thats the answer

OpenStudy (anonymous):

no its not. its not an option.

OpenStudy (amistre64):

the proof is pretty simple xcrypt suppose we have an arith seq: a1 to an

OpenStudy (amistre64):

try doing the math to finish it out; im not your nanny

OpenStudy (anonymous):

i know its a negative.

OpenStudy (anonymous):

so is it like, -2,108?

OpenStudy (amistre64):

\[A = a_1+a_2+a_3+...+a_n\] to sum this up we can add it together twice, but in a useful way, reverse the second set \[A = a_1+a_2+a_3+...+a_n\]\[A = a_n+a_{n-1}+a_{n-2}+...+a_1\] that gets us 2a = (a1+an) + (a1+an) + ... + (a1+an) ; n times

OpenStudy (anonymous):

was my answer correct?

OpenStudy (anonymous):

@Ethan96 use a calculator? Or just type in the equation in google

OpenStudy (amistre64):

17*(-126) i believe

OpenStudy (anonymous):

-2,142

OpenStudy (amistre64):

divide both sides by 2 to get: \[A=\frac{n}{2}(a_1+a_n)\]

OpenStudy (anonymous):

okay,so -2,142 is correct? :D

OpenStudy (amistre64):

that looks better yes

OpenStudy (anonymous):

THANKS

OpenStudy (amistre64):

theres other proof, but that ones the one i like

OpenStudy (anonymous):

hmm great, thanks! What's this topic called in math? Never seen it before

OpenStudy (amistre64):

series and sequences is usually when it comes about, but that little tidbit also comes up in discrete math

OpenStudy (anonymous):

okay thanks ^^